OpenStudy (algtrigcalc):

X^2+y^2=2(2y-x+2) How do I tell this equation is a parabola, circle, ellipse, or hyperbola without graphing?

satellite73 (satellite73):

look to see if the coefficients on the \(x^2\) and \(y^2\) terms are equal or not

satellite73 (satellite73):

you don't even have to multiply on the right or collect like terms or anything to see it

OpenStudy (algtrigcalc):

So I only have to see if x^2 and y^2 and equal for it to be a circle?

satellite73 (satellite73):

maybe i was not clear the "coefficients" have to be equal \[2x^2+3y^2+\text{stuff}=\text{whatever}\] is an ellipse

satellite73 (satellite73):

\[2x^2-3y^2+\text{stuff}=\text{whatever}\] is a hyperbola

OpenStudy (sunnnystrong):

Circles are going to be in the form ax^2+by^2+cx+dx+e Where a b c d e are any real number

satellite73 (satellite73):

\[x^2+y^2+\text{stuff}=\text{whatever}\] is a circle as is \[3x^2+3y^2+\text{stuff}=\text{whatever}\] because the coefficients are the same

satellite73 (satellite73):

@sunnnystrong actually no, for a circle the coefficients for the square terms must be equal

OpenStudy (algtrigcalc):

Thanks

OpenStudy (sunnnystrong):

Okay a b c d e are any real number but the squaredterms must be equal. Exucse im using openstudy mobile and its terrible